Documenta Mathematica

Band 6, 2001

Raymond Brummelhuis, Norbert Röhrl, Heinz Siedentop Stability of the Relativistic Electron-Positron Field of Atoms in Hartree-Fock Approximation:

Heavy Elements  1–9

Matei Toma Compact Moduli Spaces of Stable Sheaves

over Non-Algebraic Surfaces  11–29

Jörg Winkelmann How Frequent Are Discrete Cyclic Subgroups

of Semisimple Lie Groups?  31–37

Manfred Streit and Jürgen Wolfart

Cyclic Projective Planes and Wada Dessins  39–68

C. Deninger

On the Γ-Factors of Motives II  69–97

R. Skip Garibaldi, Anne Quéguiner-Mathieu, Jean-Pierre Tignol Involutions and Trace Forms on Exterior Powers

of a Central Simple Algebra  99–120

Henning Krause On Neeman’s Well Generated

Triangulated Categories  121–126

Jérôme Chabert and Siegfried Echterhoff Permanence Properties

of the Baum-Connes Conjecture  127–183

Jan Stevens Rolling Factors Deformations

and Extensions of Canonical Curves  185–226

Daniel Huybrechts Products of Harmonic Forms

and Rational Curves  227–239

Mikael Rørdam

Extensions of Stable C^{∗}-Algebras  241–246

Brooks Roberts

Global L-Packets for GSp(2) and Theta Lifts  247–314

Annette Werner Compactification of the Bruhat-Tits Building of PGL

by Lattices of Smaller Rank  315–341

Michael Anshelevich Partition-Dependent Stochastic Measures

and q-Deformed Cumulants  343–384

Oleg H. Izhboldin and Ina Kersten

Excellent Special Orthogonal Groups  385–412

Jörg Winkelmann Realizing Countable Groups As Automorphism Groups

of Riemann Surfaces  413–417

Rutger Noot Lifting Galois Representations,

and a Conjecture of Fontaine and Mazur  419–445

P. Schneider, J. Teitelbaum

p-Adic Fourier Theory  447–481

Amnon Neeman On the Derived Category

of Sheaves on a Manifold  483–488

Grégory Berhuy, David B. Leep Divisible Subgroups of Brauer Groups

and Trace Forms of Central Simple Algebras  489–500

D. Burns and M. Flach Tamagawa Numbers for Motives with

(Non-Commutative) Coefficients  501–570

Jürgen Hausen A Generalization of Mumford’s

Geometric Invariant Theory  571–592